# Higher-Order Regularization Learning on Hypergraphs

### v1

@@ 4.2.2 Spectral convergence of $\mathcal{L}_{n}^{(q)}$
We define $S_{n}\in\Sigma_{n,i-1}$ to be the span of the orthonormal set $(\psi_{n,1},\dots,\psi_{n,i-1})$. By the [16, Max-min theorem], we have
@@
By the induction hypothesis, for each $\ell=1,\dots,i-1$, we have the convergence of eigenvalues and hence
$\lim_{n\to\infty}\beta_{n,i}=\lim_{n\to\infty}\langle\mathcal{L}_{n}^{(q)}\psi_{n,\ell},\psi_{n,\ell}\rangle_{\mathrm{L}^{2}(\mu_{n})}=\beta_{i}<\infty.$
This uniform boundedness implies that $\sup_{n}\langle\mathcal{L}_{n}^{(q)}\psi_{n,\ell},\psi_{n,\ell}\rangle_{\mathrm{L}^{2}(\mu_{n})}<\infty$ for $1\leq\ell\leq i-1$. We use the same compactness argument as above and a diagonal argument to obtain subsequences - which, to lighten notation, we do not relabel - $(\mu_{n},\psi_{n,\ell})$ converging to $(\mu,h_{\ell})$ in $\mathrm{TL}^{2}$ for some $h_{\ell}\in\mathrm{L}^{2}(\mu)$. Moreover, by [36, Proposition 2.6] and recalling that $\langle\psi_{n,\ell},\psi_{n,j}\rangle_{\mathrm{L}^{2}(\mu_{n})}=0$, we have orthonormality in the limit
@@
By the $\limsup$-inequality in [85, Theorem 3.5], we obtain $w_{n}\in\mathrm{L}^{2}(\mu_{n})$ such that $(\mu_{n},w_{n})\to(\mu,w)$ in $\mathrm{TL}^{2}(\Omega)$ and $\limsup_{n\to\infty}\langle\mathcal{L}_{n}^{(q)}w_{n},w_{n}\rangle_{\mathrm{L}^{2}(\mu_{n})}\leq\langle w,\mathcal{L}^{(q)}w\rangle_{\mathrm{L}^{2}(\mu)}.$ Let us define the projection of $w_{n}$ onto the orthogonal complement of $S_{n}$ as
@@
By construction, $\tilde{w}_{n}\in{S_{n}}^{\perp}$. Moreover, from [36, Proposition 2.6], we have $\langle w_{n},\psi_{n,\ell}\rangle_{\mathrm{L}^{2}(\mu_{n})}\to\langle w,h_{\ell}\rangle_{\mathrm{L}^{2}(\mu)}=0$ (since $w\in S^{\perp}$) as $n\to\infty$ for all $1\leq\ell\leq i-1$, and hence, it is straight-forward to check that $(\mu_{n},\tilde{w}_{n})\to(\mu,w)$ in $\mathrm{TL}^{2}(\Omega)$.
@@
(58) $\limsup_{n\to\infty}\langle\mathcal{L}_{n}^{(q)}\tilde{w}_{n},\tilde{w}_{n}\rangle_{\mathrm{L}^{2}(\mu_{n})}\leq\limsup_{n\to\infty}\langle\mathcal{L}_{n}^{(q)}w_{n},w_{n}\rangle_{\mathrm{L}^{2}(\mu_{n})}-\sum_{\ell=1}^{i-1}\beta_{n,\ell}\langle w_{n},\psi_{n,\ell}\rangle_{\mathrm{L}^{2}(\mu_{n})}^{2}\leq\langle w,\mathcal{L}^{(q)}w\rangle_{\mathrm{L}^{2}(\mu)}$
where we used the $\limsup$-inequality of [85, Theorem 3.5] and the fact that $\langle w_{n},\psi_{n,\ell}\rangle_{\mathrm{L}^{2}(\mu_{n})}\to 0$ for (58).
Since $(\mu_{n},\tilde{w}_{n})\to(\mu,w)$ in $\mathrm{TL}^{2}(\Omega)$ and $\|w\|_{\mathrm{L}^{2}(\mu)}=1$, [36, Proposition 2.6] implies $\lim_{n\to\infty}\|\tilde{w}_{n}\|_{\mathrm{L}^{2}(\mu_{n})}=1$ and we can thus define
@@
$\lim_{n\to\infty}\beta_{n,i}$ $=\lim_{n\to\infty}\min_{v\in S_{n}^{\perp},\,\|v\|_{\mathrm{L}^{2}(\mu_{n})}}\langle\mathcal{L}_{n}^{(q)}v,v\rangle_{\mathrm{L}^{2}(\mu_{n})}$
@@
Base case $i=1$
We covered the convergence of $\psi_{n,1}$ to $\psi_{1}$ in $\mathrm{TL}^{2}(\Omega)$ in the base case of the convergence of eigenvalues. Regarding the projections, assume that $(\mu_{n},v_{n})\to(\mu,v)$ in $\mathrm{TL}^{2}(\Omega)$. Since by Assumption S.1 $\Omega$ is connected, the first eigenvalue $\overline{\gamma}_{1}=0$ is simple, and $\mathrm{Proj}_{1}(v)$ corresponds to the constant function equal to the mean of $v$ with respect to $\mu$, that is, $\mathrm{Proj}_{1}(v)=\langle v,\mathds{1}\rangle_{\mathrm{L}^{2}(\mu)}.$ Similarly, convergence of eigenvalue multiplicities (62) implies that for $n$ large enough, $E_{n,i}$ is one-dimensional. In this case, $\mathrm{Proj}_{n,1}(v_{n})$ is the constant vector equal to $\langle v_{n},\mathbf{1}\rangle_{\mathrm{L}^{2}(\mu_{n})}$. By [36, Proposition 2.6], we have
@@
By [36, Proposition 2.6] and the fact that $\|\psi_{n,j}\|_{\mathrm{L}^{2}(\mu_{n})}=1$, we also have $\|h_{j}\|_{\mathrm{L}^{2}(\mu)}=1$, implying that
@@
By using the convergence of eigenvalues, the $\liminf$-inequality of [85, Theorem 3.5] and (63), we obtain
(64) $\bar{\gamma}_{i}=\gamma_{j}=\lim_{n\to\infty}\beta_{n,j}=\liminf_{n\to\infty}\langle\mathcal{L}_{n}^{(q)}\psi_{n,j},\psi_{n,j}\rangle_{\mathrm{L}^{2}(\mu_{n})}\geq\langle\mathcal{L}^{(q)}h_{j},h_{j}\rangle_{\mathrm{L}^{2}(\mu)}\geq\bar{\gamma}_{i}.$
This implies that $\langle\mathcal{L}^{(q)}h_{j},h_{j}\rangle_{\mathrm{L}^{2}(\mu)}=\bar{\gamma}_{i}$ and, (63) also allows us to deduce that $\mathrm{Proj}_{r}(h_{j})=0$ for all $r\neq i$. We conclude that $h_{j}$ is an eigenvector of $\mathcal{L}^{(q)}$ with eigenvalue $\bar{\gamma}_{i}$, establishing the convergence of eigenvectors.
@@
Similarly to the above, for each $j=1,\dots,s(i)$, the sequence $\{v_{n,j}\}_{n\in\mathbb{N}}$ is precompact in $\mathrm{TL}^{2}$ and - without relabeling the subsequences - we may assume that $(\mu_{n},v_{n,j})\to(\mu,v_{j})$ in $\mathrm{TL}^{2}(\Omega)$ for some $v_{j}\in\mathrm{L}^{2}(\mu)$. From [36, Proposition 2.6], it follows that each $v_{j}$ satisfies $\|v_{j}\|_{\mathrm{L}^{2}(\mu_{j})}=1$, and that the family $\{v_{1},\dots,v_{s(i)}\}$ is orthonormal with respect to $\langle\cdot,\cdot\rangle_{\mathrm{L}^{2}(\mu)}$. Moreover, by the convergence of eigenvectors, each $v_{j}$ lies in the limiting eigenspace $E_{i}$, so that $\{v_{1},\dots,v_{s(i)}\}$ forms an orthonormal basis for $E_{i}$. Hence, the projection $\mathrm{Proj}_{i}$ can be written for $v\in\mathrm{L}^{2}(\mu)$ as
@@
Now, since $(\mu_{n},w_{n})\to(\mu,w)$ and $(\mu_{n},v_{n,j})\to(\mu,v_{j})$ in $\mathrm{TL}^{2}(\Omega)$, we apply [36, Proposition 2.6] to conclude. ∎
@@
We want to proceed as in the proof of [22, Theorem 2] where the analogous statement is proven for $\Delta_{n,\varepsilon_{n}}^{s}$ (see also the proof of Proposition 4.5 for a similar argument). In particular, the authors mainly rely on the fact that the eigenpairs of $\Delta_{n,\varepsilon_{n}}$ converge to the eigenpairs of $\Delta_{\rho}$. In our case, by Proposition 4.7, the eigenpairs of $\mathcal{L}_{n}^{(q)}$ converge to eigenpairs of $\mathcal{L}^{(q)}$ and we can therefore apply the same argument to conclude. ∎
@@
We want to proceed as in the proof of [87, Lemma 4.19] where the results was shown for the energy $\langle v_{n},\Delta_{n,\varepsilon_{n}}v_{n}\rangle_{\mathrm{L}^{2}(\mu_{n})}$. In particular, the proof relies on the following elements:
@@
    There exists a constant $C(u)$ such that $\|\Delta_{\rho}(u)-\Delta_{n,\varepsilon_{n}}(u)\|_{\mathrm{L}^{2}(\mu_{n})}\leq C(u)\varepsilon_{n}\to 0$ [34, Theorem 2.8].
@@
    the fact that $\|\mathcal{L}_{n}^{(q)}(u)-\mathcal{L}^{(q)}(u)\|_{\mathrm{L}^{2}(\mu)}\leq C(u)\sum_{k=1}^{q}\lambda_{k}\varepsilon_{n}^{(k)}\to 0$. Specifically, let $E_{k}$ be the set such that [34, Theorem 2.8] holds for $\varepsilon_{n}^{(k)}$: we can apply the latter result since the assumptions that $p_{1}\leq\cdots\leq p_{q}$ and $\varepsilon_{n}^{(1)}>\cdots\varepsilon_{n}^{(q)}$ imply that (65) holds for any $1\leq k\leq q$. We know from the proof of Proposition 4.4 that, for any $\alpha>1$, there exists $0<c<C$ and $\varepsilon_{0}>0$ such that $\mathbb{P}\left(\cap_{k=1}^{q}E_{k}\right)\geq 1-Cn^{-\alpha}-Cne^{-cn\left(\varepsilon_{n}^{(q)}\right)^{d+4p_{q}}}$ as long as $\varepsilon_{0}\geq\varepsilon_{n}^{(1)}>\dots>\varepsilon_{n}^{(q)}$. On this intersection, we have
@@
    where we used [31, Theorem 2.8] for the inequality. The last term tends to 0 and, by applying the Borel-Cantelli lemma with (65), we can show that this convergence holds $\mathbb{P}$-a.e..
We therefore apply the same argument as in [87, Lemma 4.19] to deduce the claim. ∎
@@
We are going to proceed as in the proof of Proposition 4.5 where the same result is proven for the truncated energy of a single Laplacian matrix $\Delta_{n,\varepsilon_{n}}^{s}$. If we replace the latter by $\mathcal{L}_{n}^{(q)}$, [87, Lemma 4.19] by Proposition 4.9, the convergence of eigenpairs by Proposition 4.7 and [87, Proposition 4.21] by Corollary 4.8 the same proof applies. ∎

### v2

@@ 4.2.2 Spectral convergence of $\mathcal{L}_{n}^{(q)}$
We define $S_{n}\in\Sigma_{n,i-1}$ to be the span of the orthonormal set $(\psi_{n,1},\dots,\psi_{n,i-1})$. By [Chavel, Max-min theorem], we have
@@
By the induction hypothesis, for each $1\leq\ell\leq i-1$, we have the convergence of eigenvalues and hence
$\lim_{n\to\infty}\beta_{n,\ell}=\lim_{n\to\infty}\langle\mathcal{L}_{n}^{(q)}\psi_{n,\ell},\psi_{n,\ell}\rangle_{\mathrm{L}^{2}(\mu_{n})}=\beta_{\ell}<\infty.$
This uniform boundedness implies that $\sup_{n}\langle\mathcal{L}_{n}^{(q)}\psi_{n,\ell},\psi_{n,\ell}\rangle_{\mathrm{L}^{2}(\mu_{n})}<\infty$ for $1\leq\ell\leq i-1$. We use the same compactness argument as above and a diagonal argument to obtain subsequences - which, to lighten notation, we do not relabel - $(\mu_{n},\psi_{n,\ell})$ converging to $(\mu,h_{\ell})$ in $\mathrm{TL}^{2}$ for some $h_{\ell}\in\mathrm{L}^{2}(\mu)$. Moreover, by [GARCIATRILLOS2018239, Proposition 2.6] and recalling that $\langle\psi_{n,\ell},\psi_{n,j}\rangle_{\mathrm{L}^{2}(\mu_{n})}=0$, we have orthonormality in the limit
@@
By the $\limsup$-inequality in [weihs2025Hypergraphs, Theorem 3.5], we obtain $w_{n}\in\mathrm{L}^{2}(\mu_{n})$ such that $(\mu_{n},w_{n})\to(\mu,w)$ in $\mathrm{TL}^{2}(\Omega)$ and $\limsup_{n\to\infty}\langle\mathcal{L}_{n}^{(q)}w_{n},w_{n}\rangle_{\mathrm{L}^{2}(\mu_{n})}\leq\langle w,\mathcal{L}^{(q)}w\rangle_{\mathrm{L}^{2}(\mu)}.$ Let us define the projection of $w_{n}$ onto the orthogonal complement of $S_{n}$ as
@@
By construction, $\tilde{w}_{n}\in{S_{n}}^{\perp}$. Moreover, from [GARCIATRILLOS2018239, Proposition 2.6], we have $\langle w_{n},\psi_{n,\ell}\rangle_{\mathrm{L}^{2}(\mu_{n})}\to\langle w,h_{\ell}\rangle_{\mathrm{L}^{2}(\mu)}=0$ (since $w\in S^{\perp}$) as $n\to\infty$ for all $1\leq\ell\leq i-1$, and hence, it is straight-forward to check that $(\mu_{n},\tilde{w}_{n})\to(\mu,w)$ in $\mathrm{TL}^{2}(\Omega)$.
@@
$\limsup_{n\to\infty}\langle\mathcal{L}_{n}^{(q)}\tilde{w}_{n},\tilde{w}_{n}\rangle_{\mathrm{L}^{2}(\mu_{n})}$ $\leq\limsup_{n\to\infty}\langle\mathcal{L}_{n}^{(q)}w_{n},w_{n}\rangle_{\mathrm{L}^{2}(\mu_{n})}-\sum_{\ell=1}^{i-1}\beta_{n,\ell}\langle w_{n},\psi_{n,\ell}\rangle_{\mathrm{L}^{2}(\mu_{n})}^{2}$
(59) $\leq\langle w,\mathcal{L}^{(q)}w\rangle_{\mathrm{L}^{2}(\mu)}$
where we used the $\limsup$-inequality of [weihs2025Hypergraphs, Theorem 3.5] and the fact that $\beta_{n,\ell}\langle w_{n},\psi_{n,\ell}\rangle_{\mathrm{L}^{2}(\mu_{n})}^{2}\geq 0$ for (59).
Since $(\mu_{n},\tilde{w}_{n})\to(\mu,w)$ in $\mathrm{TL}^{2}(\Omega)$ and $\|w\|_{\mathrm{L}^{2}(\mu)}=1$, [GARCIATRILLOS2018239, Proposition 2.6] implies $\lim_{n\to\infty}\|\tilde{w}_{n}\|_{\mathrm{L}^{2}(\mu_{n})}=1$ and we can thus define
@@
$\lim_{n\to\infty}\beta_{n,i}$ $=\lim_{n\to\infty}\min_{v\in S_{n}^{\perp},\,\|v\|_{\mathrm{L}^{2}(\mu_{n})=1}}\langle\mathcal{L}_{n}^{(q)}v,v\rangle_{\mathrm{L}^{2}(\mu_{n})}$
@@
Base case $i=1$.
We covered the convergence of $\psi_{n,1}$ to $\psi_{1}$ in $\mathrm{TL}^{2}(\Omega)$ in the base case of the convergence of eigenvalues. Regarding the projections, assume that $(\mu_{n},v_{n})\to(\mu,v)$ in $\mathrm{TL}^{2}(\Omega)$. Since by Assumption S.1 $\Omega$ is connected, the first eigenvalue $\overline{\gamma}_{1}=0$ is simple, and $\mathrm{Proj}_{1}(v)$ corresponds to the constant function equal to the mean of $v$ with respect to $\mu$, that is, $\mathrm{Proj}_{1}(v)=\langle v,\mathds{1}\rangle_{\mathrm{L}^{2}(\mu)}.$ Similarly, convergence of eigenvalue multiplicities (63) implies that for $n$ large enough, $E_{n,i}$ is one-dimensional. In this case, $\mathrm{Proj}_{n,1}(v_{n})$ is the constant vector equal to $\langle v_{n},\mathbf{1}\rangle_{\mathrm{L}^{2}(\mu_{n})}$. By [GARCIATRILLOS2018239, Proposition 2.6], we have
@@
By [GARCIATRILLOS2018239, Proposition 2.6] and the fact that $\|\psi_{n,j}\|_{\mathrm{L}^{2}(\mu_{n})}=1$, we also have $\|h_{j}\|_{\mathrm{L}^{2}(\mu)}=1$, implying that
@@
By using the convergence of eigenvalues, the $\liminf$-inequality of [weihs2025Hypergraphs, Theorem 3.5] and (64), we obtain
$\bar{\gamma}_{i}=\gamma_{j}=\lim_{n\to\infty}\beta_{n,j}=\liminf_{n\to\infty}\langle\mathcal{L}_{n}^{(q)}\psi_{n,j},\psi_{n,j}\rangle_{\mathrm{L}^{2}(\mu_{n})}\geq\langle\mathcal{L}^{(q)}h_{j},h_{j}\rangle_{\mathrm{L}^{2}(\mu)}\geq\bar{\gamma}_{i}.$
This implies that $\langle\mathcal{L}^{(q)}h_{j},h_{j}\rangle_{\mathrm{L}^{2}(\mu)}=\bar{\gamma}_{i}$ and (64) also allows us to deduce that $\mathrm{Proj}_{r}(h_{j})=0$ for all $r\neq i$. Indeed, suppose that $\|\mathrm{Proj}_{r}(h_{j})\|_{\mathrm{L}^{2}(\mu)}>0$ for $r>i$. Then (following the same spectral decomposition preceding (64)), we have $\langle\mathcal{L}^{(q)}h_{j},h_{j}\rangle_{\mathrm{L}^{2}(\mu)}>\bar{\gamma}_{i}$ which is a contradiction. We conclude that $h_{j}$ is an eigenvector of $\mathcal{L}^{(q)}$ with eigenvalue $\bar{\gamma}_{i}$, establishing the convergence of eigenvectors.
@@
Similarly to the above, for each $j=1,\dots,s(i)$, the sequence $\{v_{n,j}\}_{n\in\mathbb{N}}$ is precompact in $\mathrm{TL}^{2}$ and - without relabeling the subsequences - we may assume that $(\mu_{n},v_{n,j})\to(\mu,v_{j})$ in $\mathrm{TL}^{2}(\Omega)$ for some $v_{j}\in\mathrm{L}^{2}(\mu)$. From [GARCIATRILLOS2018239, Proposition 2.6], it follows that each $v_{j}$ satisfies $\|v_{j}\|_{\mathrm{L}^{2}(\mu_{j})}=1$, and that the family $\{v_{1},\dots,v_{s(i)}\}$ is orthonormal with respect to $\langle\cdot,\cdot\rangle_{\mathrm{L}^{2}(\mu)}$. Moreover, by the convergence of eigenvectors, each $v_{j}$ lies in the limiting eigenspace $E_{i}$, so that $\{v_{1},\dots,v_{s(i)}\}$ forms an orthonormal basis for $E_{i}$. Hence, the projection $\mathrm{Proj}_{i}$ can be written for $v\in\mathrm{L}^{2}(\mu)$ as
@@
Now, since $(\mu_{n},w_{n})\to(\mu,w)$ and $(\mu_{n},v_{n,j})\to(\mu,v_{j})$ in $\mathrm{TL}^{2}(\Omega)$, we apply [GARCIATRILLOS2018239, Proposition 2.6] to conclude. ∎
@@
We want to proceed as in the proof of [Stuart, Theorem 2] where the analogous statement is proven for $\Delta_{n,\varepsilon_{n}}^{s}$ (see also the proof of Proposition 4.5 for a similar argument). In particular, the authors mainly rely on the fact that the eigenpairs of $\Delta_{n,\varepsilon_{n}}$ converge to the eigenpairs of $\Delta_{\rho}$. In our case, by Proposition 4.7, the eigenpairs of $\mathcal{L}_{n}^{(q)}$ converge to eigenpairs of $\mathcal{L}^{(q)}$ and we can therefore apply the same argument to conclude. ∎
@@
We want to proceed as in the proof of [weihs2023consistency, Lemma 4.19] where the results was shown for the energy $\langle v_{n},\Delta_{n,\varepsilon_{n}}v_{n}\rangle_{\mathrm{L}^{2}(\mu_{n})}$. In particular, the proof relies on the following elements:
@@
    There exists a constant $C(u)$ such that $\|\Delta_{\rho}(u)-\Delta_{n,\varepsilon_{n}}(u)\|_{\mathrm{L}^{2}(\mu_{n})}\leq C(u)\varepsilon_{n}\to 0$ [trillos2022rates, Theorem 2.8].
@@
    the fact that $\|\mathcal{L}_{n}^{(q)}(u)-\mathcal{L}^{(q)}(u)\|_{\mathrm{L}^{2}(\mu)}\leq C(u)\sum_{k=1}^{q}\lambda_{k}\varepsilon_{n}^{(k)}\to 0$. Specifically, let $E_{k}$ be the set such that [trillos2022rates, Theorem 2.8] holds for $\varepsilon_{n}^{(k)}$: we can apply the latter result since the assumptions that $p_{1}\leq\cdots\leq p_{q}$ and $\varepsilon_{n}^{(1)}>\cdots\varepsilon_{n}^{(q)}$ imply that (65) holds for any $1\leq k\leq q$. We know from the proof of Proposition 4.4 that, for any $\alpha>1$, there exists $0<c<C$ and $\varepsilon_{0}>0$ such that $\mathbb{P}\left(\cap_{k=1}^{q}E_{k}\right)\geq 1-Cn^{-\alpha}-Cne^{-cn\left(\varepsilon_{n}^{(q)}\right)^{d+4p_{q}}}$ as long as $\varepsilon_{0}\geq\varepsilon_{n}^{(1)}>\dots>\varepsilon_{n}^{(q)}$. On this intersection, we have
@@
    where we used [Trillos, Theorem 2.8] for the inequality. The last term tends to 0 and, by applying the Borel-Cantelli lemma with (65), we can show that this convergence holds $\mathbb{P}$-a.e..
We therefore apply the same argument as in [weihs2023consistency, Lemma 4.19] to deduce the claim. ∎
@@
We are going to proceed as in the proof of Proposition 4.5 where the same result is proven for the truncated energy of a single Laplacian matrix $\Delta_{n,\varepsilon_{n}}^{s}$. If we replace the latter by $\mathcal{L}_{n}^{(q)}$, [weihs2023consistency, Lemma 4.19] by Proposition 4.9, the convergence of eigenpairs by Proposition 4.7 and [weihs2023consistency, Proposition 4.21] by Corollary 4.8 the same proof applies. ∎